In unknown volume of 40% w/w solutions of h2so4 of density 1.4g/ml is diluted until it`s density becomes 1.05per ml . calculate %w/w& %w/v

Grade:11

1 Answers

Vikas TU
6 years ago
Dear Student,
gram atomic weight of h2so4= 98 
40% of weight of the solution= 98; 
Gram Weight of solution= 245. 
After weakening, volume= 245/1.05=233.33ml. 
In this manner, w/v= 98/233.33= 42% 
w/w% wtill= 40% 
Cheers!!
Regards,
Vikas (B. Tech. 4th year
Thapar University)